Punit gets 10% less marks than Bini. What percentage of marks which Bini gets more than punit?
1
Expert's answer
2012-06-07T08:05:49-0400
Punit gets 10% less marks than Bini. What percentage of marks which Bini gets more than punit?

Let Bini has B marks. Then Punit has P = (1-0.1)B = 0.9B marks.

B = (1/0.9)P = 1.(1)P.

So, Bini has 11,(1)% more marks than Punit.